Alberto made a trip to his friends house and back. The trip there took three hours and the trip back took four hours. He averaged 9.8 mph faster on the trip there than on the return trip. Find Alberto's average speed on the outbound trip.
Answer by mananth(16946)

let the speed on return be x mph : time = 4 hours
speed going there = x+9.8 mph : Time = 3 hours
Distance is the same forward and return
d=rt
4x= 3(x+9.8)
4x= 3x+29.4
x= 29.4 mph